Topic is over effective rate of interest.

If A(4)=1000 and i_n=0.01n, find A(7)
Answer #1

Let initial amount be P

A(4) = P*(1+i1)(1+i2)(1+i3)(1+i4)

=> 1000 = P*(1+0.01*1)(1+0.01*2)(1+0.01*3)(1+0.01*4)

=> P = $906.17

Hence, A(7) = P(1+i1)(1+i2)(1+i3)(1+i4)(1+i5)(1+i6)(1+i7)

= 906.17*(1+0.01*1)(1+0.01*2)(1+0.01*3)(1+0.01*4)(1+0.01*5)(1+0.01*6)(1+0.01*7)

= $1190.91